OpenStudy (anonymous):

scientific notation simplify 52.9/10^-7

OpenStudy (ash2326):

@Diana97 do you know what's \[\frac{1}{a^{-5}}=??\]

OpenStudy (anonymous):

no

OpenStudy (ash2326):

If I bring the denominator to the numerator the power will change its sign\ so \[\frac{1}{a^{-5}}=a^5\]

OpenStudy (ash2326):

@Diana97 do you get this part?

OpenStudy (anonymous):

yea

OpenStudy (ash2326):

we have \[52.9/10^{-7}\] bring the denominator to numerator, would you?

OpenStudy (anonymous):

i understand a little bet now i want to check my answers